find the perimeter of a rectangle ABCD with a side of 8 inches and a diagonal of 10 inches?
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16949)

find the perimeter of a rectangle ABCD with a side of 8 inches and a diagonal of 10 inches?
Let the other side be x
Side = 8 inches
Diagonal = 10 inches
10^2 = 8^2 + x^2 pythagoras
100-64=x^2
36=x^2
X=6
Perimeter = 2*(l+b)
=2(6+8)
=28 inches
mananth@hotmail.com

Answer by richwmiller(17219)

This a multiple of a 3-4-5 triangle (6-8-10)
so the other side is 6 inches
2*6+2*8=P
12+16=28 inches